OpenStudy (anonymous):

Simplify the complex fraction. 4/(x+3)/(1/x+3) A. 12x+4/x^2+3x B. 4x/3x+9 C. 4x/3x^2+10x+3 D. None of these

OpenStudy (unklerhaukus):

more brackets are required to avoid ambiguity

OpenStudy (anonymous):

\[\frac{\frac{4}{x+3}}{\frac{1}{x+3}}=\frac{4}{x+3}\times \frac{x+3}{1}=4\]

OpenStudy (anonymous):

so it's D?

OpenStudy (anonymous):

\[4/(x+3)/(1/x+3)=\frac{4}{x+3}/\left(\frac{1}{x}+3\right)=\frac{4}{\left(\frac{1}{x}+3\right) (x+3)}=\frac{4 x}{3 x^2+10 x+3} \]

OpenStudy (anonymous):

C:

OpenStudy (unklerhaukus):

@jaydemarie4, note that satellite73, and robtobey got different answers because the question as written is ambiguous

OpenStudy (anonymous):

Not in this case because of the precedence rules. The first division sign from the left is applied first. Mathematica stands by it's solution.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

You're welcome.